OpenStudy (anonymous):

Need help ASAP!!!! Write with positive exponents and then simplify. 729 to the -4/3 <--exponent!!!!!! Please helppp

OpenStudy (anonymous):

take the cubed root of 729 first

OpenStudy (anonymous):

\(729^{-\frac{4}{3}}\) the exponent has 3 parts a numerator, which is the power a denominator, which is the root a minus sign, which means take the reciprocal i.e. \[\frac{1}{\sqrt[3]{729^4}}\]

OpenStudy (anonymous):

the cubed root of 729 is 9 because \(9^3=729\)

OpenStudy (anonymous):

giving you \[\frac{1}{9^4}\]

OpenStudy (anonymous):

or if you prefer \[\frac{1}{6561}\]